Write an equation of the line that is parallel to the given line
y = 2x + 3 and the given point (-4,1).

Answer by robertb(5830)

The line must have the form y = 2x + c, where c is unknown. To find c, use the coordinates of the given point. 1 = 2*-4 + c, or 1 = -8 + c, or c = 9. Therefore the equation of the line is y = 2x + 9.